# QDM: Quadtree-Based Region-Adaptive Sparse Diffusion Models for Efficient Image Super-Resolution

### v1

@@ 1 Introduction
Super-resolution (SR) seeks to reconstruct high-resolution (HR) images from low-resolution (LR) inputs, an inherently ill-posed inverse problem due to the unknown and complex image formation processes in real-world scenarios. Existing approaches mitigate this challenge by designing empirical image formation pipelines and leveraging generative adversarial networks (GANs) [8] to approximate the LR-to-HR mapping [45, 35, 17]. Recent advances in diffusion models [10, 29, 30] have demonstrated superior image generation capabilities compared to GANs [5], positioning them as promising tools for SR. Contemporary studies achieve this either by adapting diffusion frameworks for SR tasks [27, 25, 44, 36] or exploiting pretrained text-to-image (T2I) models like Stable Diffusion [25, 23] as priors to regularize the ill-posed reconstruction [19, 38, 33, 39, 42, 40].
@@
Figure 1: The original high-resolution (HR) natural images collected from DIV2K [1] and medical images from SegRap2023 [21], are downsampled to lower resolutions. The difference heatmaps reveal that only sparse detail-rich regions—such as edges and textures—require significant refinement, while large homogeneous areas remain largely unaffected.
@@
For the first challenge, we employ a quadtree decomposition strategy [28]. This approach recursively subdivides regions until they reach a predefined homogeneity threshold, resulting in variable-sized homogeneous regions at the leaf nodes. We can then leverage this quadtree prior to refine regions only where necessary, thus reducing redundant computations. For the second challenge, existing diffusion-based SR methods [10, 30, 25, 44] typically rely on architectures, such as U-Net [26] or Transformer-based models [22], which uniformly process all spatial regions without adaptivity. We instead propose a dual-stream architecture guided by the quadtree mask. Our design processes coarse-grained patches in upstream blocks to capture global context, while dynamically routing detail-rich regions to downstream blocks for enhancing fine-grained patches. Our approach provides two main advantages: (1) it localizes computationally intensive operations to detail-rich regions, and (2) it partitions selected tokens into parallelizable chunks, eliminating the need for full-feature-map storage.
@@ 2 Related Works
Diffusion-based Image Super-Resolution. Diffusion-based models for super-resolution (SR) have been widely explored and can be roughly categorized into two approaches: train-from-scratch and pretrained-prior-based methods. Train-from-scratch techniques, initiated by SR3 [27], pioneered the application of diffusion models to SR. LDM [25] further enhanced computational efficiency by operating in a latent space, while ResShift [44] further streamlined the process to $15$ diffusion steps via residue shifting. These methods are easy to retrain, and inherently versatile, making them broadly applicable across diverse scenarios. Another research direction leverages the generative power of pretrained text-to-image (T2I) models like Stable Diffusion [25, 23]. Works like DiffBIR [19], SeeSR [38], and StableSR [33], along with others [39, 42, 40], employ various fine-tuning strategies to exploit diffusion priors for SR, achieving state-of-the-art performance. However, such methods rely heavily on large-scale T2I foundation models, limiting their adaptability to domains beyond natural images. Our approach falls within the train-from-scratch paradigm, prioritizing flexibility and broad applicability without using external pretrained models.
@@
Model Architecture for Diffusion Models. The rapid advancement of diffusion models has established U-Net architectures [26] as foundational to diffusion-based image synthesis. These architectures have demonstrated exceptional performance across tasks spanning denoising, super-resolution, and high-definition image generation [10, 30, 25, 44]. Early frameworks like DDPMs [10] augmented U-Nets with self-attention layers at coarse resolutions to enhance feature interaction. Later studies [11, 23] emphasize the critical role of additional self-attention in low-resolution stages for scaling to high-dimensional outputs. Concurrently, the Swin Transformer [20] introduces shifted-window self-attention, which confines attention to localized windows while preserving global context through window shifting. This approach has been successfully integrated into diffusion frameworks like ResShift [44] and SinSR [36] to balance scalability and efficiency. Recently, DiT [22] has emerged as a milestone, replacing U-Nets entirely with pure Transformer backbones, showcasing unprecedented scalability for large-scale generative tasks.
@@
Quadtree Representations. Quadtrees are hierarchical data structures with a long history in image processing [28]. Early works, such as [2, 7], applied quadtrees to semantic segmentation tasks using graphical models. However, in the deep learning era, their use has been primarily focused on reducing computation and memory costs. For instance, [12] presents a quadtree-based convolutional network for tasks involving sparse, binary input images. Similarly, [4] leverages an efficient quadtree representation of segmentation masks to reduce memory consumption with minimal accuracy loss. Quadtree attention [31] computes attention in a coarse-to-fine manner, reducing computational complexity from quadratic to linear.
@@ 3.1 Quadtree Prior—Identify Detail-Rich Regions
A quadtree is a hierarchical data structure where each non-leaf node has four child nodes [28]. This structure recursively partitions a two-dimensional space into four quadrants or regions. If all pixels within a sub-region exhibit similar values (i.e., low variance), the sub-region becomes a leaf node in the quadtree and is represented by a single aggregated value. In this work, we adopt a lossy compression strategy, considering a sub-region as a leaf node if the $l_{1}$ norm of the difference between its maximum and minimum values is below a threshold $s$. This approach maximizes compression while preserving essential image details.
Using this scheme, we construct a hierarchical quadtree structure from a low-resolution (LR) image $Y_{0}$. The bottom level of the quadtree represents the finest granularity, corresponding to the resolution of the LR image. A key assumption in our approach is that LR and HR images share the same structural hierarchy at upper levels. Consequently, leaf nodes at higher levels require minimal refinement, as they represent relatively homogeneous regions. In contrast, bottom-level nodes may require further subdivision and refinement to capture finer details. Thus, the primary objective is to efficiently identify the positions of these bottom-level leaf nodes. We propose to introduce a binary mask $M$ to store leaf node positions efficiently. Here, “1”s indicate the presence of leaf nodes at the bottom level, while “0”s correspond to homogeneous regions. These binary masks are efficiently generated using Algorithm 1, which is optimized for parallel execution on GPUs. Our method constructs each binary mask with minimal computational overhead, making it well-suited for real-time applications and large-scale data processing. Crucially, it’s important to note that the generated mask can have the same spatial resolution with latent feature maps rather than requiring alignment with the input LR image, enabling seamless compatibility with latent diffusion models [25].
@@
3:$l=\max\left(\log_{2}{\frac{H}{H_{m}}},\log_{2}{\frac{W}{W_{m}}}\right)$
4:$l=\max\left(\lceil l\rceil,1\right)$
5:$V^{+}\leftarrow$ Max pooling on $I$ with kernel size $2^{l}\times 2^{l}$
6:$V^{-}\leftarrow$ Max pooling on $-I$ with kernel size $2^{l}\times 2^{l}$
7:$M\leftarrow\left[\left(V^{+}+V^{-}\right)\geq s\right]$
8:$M\leftarrow$ Nearest-neighbor interpolation on $M$ to the size $(H_{m},W_{m})$
@@ 3.2 Mask-Guided Sparse Diffusion Process
Given a LR image $Y_{0}$, our goal is to generate the corresponding HR image $X_{0}$. Following standard notation in the diffusion literature, the subscript denotes the diffusion time step, where 00 corresponds to data and $1$ to pure noise. As discussed in Section 3.1, refining large homogeneous regions in an LR image is often unnecessary. To address this, we propose a mask-guided diffusion process, where the mask, often being sparse, is derived from the quadtree representation of the LR image. This approach reduces computational redundancy and reallocates model capacity toward refining detail-rich regions. In the following, we introduce the forward and reverse processes of QDM, along with the corresponding training objective.
Figure 2: Dual-Stream Architecture Overview. Our transformer-based model employs two complementary processing streams guided by a quadtree mask for input-adaptive computation. The upstream (top) extracts global context via Diffusion Transformer (DiT) blocks operating on patches with size 8$\times$8, while the downstream (bottom) refines patches with size 2$\times$2 exclusively in mask-selected regions. Mask $M$ dynamically controls computational overhead by routing computationally intensive processing only to detail-rich areas. Selected tokens are partitioned into parallelizable chunks, each processed independently through cross-attention blocks. Final predictions combine the coarse upstream outputs with the refined downstream results.
Forward Process. In the forward process, we gradually add noise to the HR image $Y_{0}$ conditioned on the LR image $X_{0}$. In contrast to vanilla diffusion models where the forward process adds noise everywhere, our approach only adds noise at active sites, i.e., detail-rich regions where the redundancy is low. In particular, we first obtain a mask $M$ of active sites via Algorithm 1, where entries of $M$ with value 1 indicate the active sites. Inspired by the work [44], we design the following transition distribution of the forward process that focuses solely on these active sites,
@@
Reverse Process. In the reverse process, we learn to perform denoise to recover the HR image. In particular, we have the following posterior distribution
@@
where $p(X_{T}\mid Y_{0},M)={\mathcal{N}}\left(X_{T};Y_{0},\mathrm{\text{diag}}(\kappa^{2}M)\right)$ and $p_{\bm{\theta}}\left(X_{t-1}\mid X_{t},Y_{0},M\right)$ is the reverse transition probability. Following common practice in diffusion models [29, 10, 30, 44], we parameterize the reverse transition probability as:
@@
$\hat{\bm{\mu}}=\frac{\eta_{t-1}}{\eta_{t}}M\odot X_{t}+\frac{\alpha_{t}}{\eta_{t}}M\odot f_{\bm{\theta}}(X_{t},Y_{0},t)$
$\qquad+(1-M)\odot f_{\bm{\theta}}(X_{t},Y_{0},t),$ (4)
where $f_{\bm{\theta}}$ is a deep neural network with parameters $\bm{\theta}$, aiming to predict $X_{0}$ and $f_{\bm{\theta}}$ is the predicted pixel values on active sites. During the reverse process, we iteratively denoise the active (detail-rich) regions while performing minimal deterministic refinement on non-active (homogeneous) areas to preserve their structures. This process is illustrated in Figure 3. The training objective is to minimize the following mean squared error,
@@
We also adopted the non-uniform geometric schedule for the diffusion time $t$ as proposed in [44].
Extension to Latent Space. To reduce computational costs, recent studies [25, 22, 44] propose training diffusion models in the latent space of a pre-trained variational autoencoder (VAE). In our framework, the encoder compresses images, reducing their spatial dimensions, which introduce a size mismatch between the compressed latent maps and the low-quality images. Fortunately, Alg. 1 can generate masks that align with the resolution of the latent maps. The spatial locations of active regions in the original image are preserved in the latent space due to the translation-equivariant properties of the convolutional encoder. Adapting our model to the latent space only requires replacing $X_{0}$ and $Y_{0}$ with their latent representations, along with the appropriately sized mask $M$.
@@ 3.3 Model Architecture
Fully transformer-based architectures [22] have outperformed the conventional U-Net models [26] in many tasks like image super-resolution. However, existing approaches are limited by their computationally expensive full-sequence attention mechanisms. This leads to substantial memory consumption and inefficient resource allocation, particularly in super-resolution tasks where only certain detail-rich regions necessitate complex processing, while others could be handled with simpler transformations.
@@
Upstream. The upstream branch utilizes a Diffusion Transformer (DiT) [22] with a large patch size of 8 to extract global features from the input image. This design enables coarse refinements in homogeneous regions while reducing the token count by 64×64$\times$64 × compared to pixel-level processing. By operating on coarse-grained patches, the model significantly lowers memory consumption and inference time while preserving global context. The large patch size expands the receptive field, improving efficiency in high-level semantic feature extraction. These global features are then propagated to downstream modules via cross-attention layers, enhancing spatial consistency in predictions.
Downstream. The downstream branch works in tandem with the upstream, refining key regions identified by the quadtree mask $M$ (Alg. 1). Given an input of size $H\times W\times C$, the image is first divided into $2\times 2$ patches, forming a feature grid of dimensions $\frac{H}{2}\times\frac{W}{2}\times C$. Following the Swin Transformer [20], this grid is further partitioned into non-overlapping $N\times N$ local windows, producing $\frac{HW}{4N^{2}}$ feature chunks of size $N^{2}\times C$. The quadtree mask $M$ (size $\frac{H}{2}\times\frac{W}{2}\times C$) selectively determines which chunks require computation: only windows overlapping with active sites (mask entries being $1$) are projected into a high-dimensional space. These selected features are then augmented with 2D positional embeddings to preserve spatial relationships.
Each feature chunk is processed through a series of attention-based operations. It first undergoes multi-head self-attention (MSA) to model local dependencies, followed by multi-head cross-attention (MCA) to incorporate global context from the upstream branch’s coarse features (dimension $C_{\text{up}}$, size $\frac{HW}{64}\times C_{\text{up}}$). The output is then passed through a two-layer feedforward network with GELU activation and time-conditioned via adaLN-Zero layers [22]. Residual connections are applied after each MSA, MCA, and feedforward operation. To improve efficiency, the upstream’s global features are pre-projected into key-value pairs ($K,V$) and cached for reuse across all chunks during inference, minimizing redundant computations. The final projection restores features to their original dimensionality.
@@
where $\mathcal{L}_{\text{LPIPS}}$ [46] measures perceptual similarity between prediction $f_{\bm{\theta}}(X_{t},Y_{0},t)$ and ground-truth $X_{0}$, with $\lambda_{3}$ balancing detail preservation against prediction accuracy.
Model Size. We apply a sequence of $N$ attention blocks for both upstream and downstream. Following DiT [22], we use standard transformer configurations that jointly scale up the number of attention blocks, the hidden size, and the number of attention heads. Specifically, we use three configs, QDM-S, QDM-B and QDM-L, aligned in size with the DiT variants. Appendix A provides more details of these configurations.
Figure 5: Comparison of QDM, DiT [22], U-Net [26], and a Swin Attention-enhanced U-Net variant [20] with high-resolution ($1024\times 1024$) inputs. The left panel shows peak memory usage, while the right panel depicts inference speed across evaluated models. All metrics were measured on an NVIDIA A100 80G GPU. For QDM, the evaluation was conducted under the most resource-intensive condition, using a full mask without adaptive computation and processing 64 chunks in parallel.

### v2

@@ 1 Introduction
Super-resolution (SR) seeks to reconstruct high-resolution (HR) images from low-resolution (LR) inputs, an inherently ill-posed inverse problem due to the unknown and complex image formation processes in real-world scenarios. Existing approaches mitigate this challenge by designing empirical image formation pipelines and leveraging generative adversarial networks (GANs) [goodfellow2020generative] to approximate the LR-to-HR mapping [zhang2021designing, wang2021real, liang2021swinir]. Recent advances in diffusion models [ho2020denoising, sohl2015deep, song2020score] have demonstrated superior image generation capabilities compared to GANs [dhariwal2021diffusion], positioning them as promising tools for SR. Contemporary studies achieve this either by adapting diffusion frameworks for SR tasks [saharia2022image, rombach2022high, yue2024resshift, wang2024sinsr] or exploiting pretrained text-to-image (T2I) models like Stable Diffusion [rombach2022high, podell2023sdxl] as priors to regularize the ill-posed reconstruction [lin2024diffbir, wu2024seesr, wang2024exploiting, wu2025one, yue2024arbitrary, yang2024pixel].
@@
Figure 1: The original high-resolution (HR) natural images collected from DIV2K [agustsson2017ntire] and medical images from SegRap2023 [luo2025segrap2023], are downsampled to lower resolutions. The difference heatmaps reveal that only sparse detail-rich regions—such as edges and textures—require significant refinement, while large homogeneous areas remain largely unaffected.
@@
For the first challenge, we employ a quadtree decomposition strategy [samet1984quadtree]. This approach recursively subdivides regions until they reach a predefined homogeneity threshold, resulting in variable-sized homogeneous regions at the leaf nodes. We can then leverage this quadtree prior to refining regions only where necessary, thus reducing redundant computations. For the second challenge, existing diffusion-based SR methods [ho2020denoising, song2020score, rombach2022high, yue2024resshift] typically rely on architectures, such as U-Net [ronneberger2015u] or Transformer-based models [peebles2023scalable], which uniformly process all spatial regions without adaptivity. We instead propose a dual-stream architecture guided by the quadtree mask. Our design processes coarse-grained patches in upstream blocks to capture global context, while dynamically routing detail-rich regions to downstream blocks for enhancing fine-grained patches. Our approach provides two main advantages: (1) it localizes computationally intensive operations to detail-rich regions, and (2) it partitions selected tokens into parallelizable chunks, eliminating the need for full-feature-map storage.
@@ 2 Related Works
Diffusion-based Image Super-Resolution. Diffusion-based models for super-resolution (SR) have been widely explored and can be roughly categorized into two approaches: train-from-scratch and pretrained-prior-based methods. Train-from-scratch techniques, initiated by SR3 [saharia2022image], pioneered the application of diffusion models to SR. LDM [rombach2022high] further enhanced computational efficiency by operating in a latent space, while ResShift [yue2024resshift] further streamlined the process to $15$ diffusion steps via residue shifting. These methods are easy to retrain, and inherently versatile, making them broadly applicable across diverse scenarios. Another research direction leverages the generative power of pretrained text-to-image (T2I) models like Stable Diffusion [rombach2022high, podell2023sdxl]. Works like DiffBIR [lin2024diffbir], SeeSR [wu2024seesr], and StableSR [wang2024exploiting], along with others [wu2025one, yue2024arbitrary, yang2024pixel], employ various fine-tuning strategies to exploit diffusion priors for SR, achieving state-of-the-art performance. However, such methods rely heavily on large-scale T2I foundation models, limiting their adaptability to domains beyond natural images. Our approach falls within the train-from-scratch paradigm, prioritizing flexibility and broad applicability without using external pretrained models.
@@
Model Architecture for Diffusion Models. The rapid advancement of diffusion models has established U-Net architectures [ronneberger2015u] as foundational to diffusion-based image synthesis. These architectures have demonstrated exceptional performance across tasks spanning denoising, super-resolution, and high-definition image generation [ho2020denoising, song2020score, rombach2022high, yue2024resshift]. Early frameworks like DDPMs [ho2020denoising] augmented U-Nets with self-attention layers at coarse resolutions to enhance feature interaction. Later studies [hoogeboom2023simple, podell2023sdxl] emphasize the critical role of additional self-attention in low-resolution stages for scaling to high-dimensional outputs. Concurrently, the Swin Transformer [liu2021swin] introduces shifted-window self-attention, which confines attention to localized windows while preserving global context through window shifting. This approach has been successfully integrated into diffusion frameworks like ResShift [yue2024resshift] and SinSR [wang2024sinsr] to balance scalability and efficiency. Recently, DiT [peebles2023scalable] has emerged as a milestone, replacing U-Nets entirely with pure Transformer backbones, showcasing unprecedented scalability for large-scale generative tasks.
@@
Quadtree Representations. Quadtrees are hierarchical data structures with a long history in image processing [samet1984quadtree]. Early works, such as [bouman1994multiscale, feng1998training], applied quadtrees to semantic segmentation tasks using graphical models. However, in the deep learning era, their use has been primarily focused on reducing computation and memory costs. For instance, [jayaraman2018quadtree] presents a quadtree-based convolutional network for tasks involving sparse, binary input images. Similarly, [chitta2020quadtree] leverages an efficient quadtree representation of segmentation masks to reduce memory consumption with minimal accuracy loss. Quadtree attention [tang2022quadtree] computes attention in a coarse-to-fine manner, reducing computational complexity from quadratic to linear.
@@ 3.1 Quadtree Prior—Identify Detail-Rich Regions
A quadtree is a hierarchical data structure where each non-leaf node has four child nodes [samet1984quadtree]. This structure recursively partitions a two-dimensional space into four quadrants or regions. If all pixels within a sub-region exhibit similar values (i.e., low variance), the sub-region becomes a leaf node in the quadtree and is represented by a single aggregated value. In this work, we adopt a lossy compression strategy, considering a sub-region as a leaf node if the $l_{1}$ norm of the difference between its maximum and minimum values is below a threshold $s$. This approach maximizes compression while preserving essential image details.
Using this scheme, we construct a hierarchical quadtree structure from a low-resolution (LR) image $Y_{0}$. The bottom level of the quadtree represents the finest granularity, corresponding to the resolution of the LR image. A key assumption in our approach is that LR and HR images share the same structural hierarchy at upper levels. Consequently, leaf nodes at higher levels require minimal refinement, as they represent relatively homogeneous regions. In contrast, bottom-level nodes may require further subdivision and refinement to capture finer details. Thus, the primary objective is to efficiently identify the positions of these bottom-level leaf nodes. We propose to introduce a binary mask $M$ to store leaf node positions efficiently. Here, “1”s indicate the presence of leaf nodes at the bottom level, while “0”s correspond to homogeneous regions. These binary masks are efficiently generated using Algorithm 1, which is optimized for parallel execution on GPUs. Our method constructs each binary mask with minimal computational overhead, making it well-suited for real-time applications and large-scale data processing. Crucially, it’s important to note that the generated mask can have the same spatial resolution with latent feature maps rather than requiring alignment with the input LR image, enabling seamless compatibility with latent diffusion models [rombach2022high].
@@
[…]
@@ 3.2 Mask-Guided Sparse Diffusion Process
Given a LR image $Y_{0}$, our goal is to generate the corresponding HR image $X_{0}$. Following standard notation in the diffusion literature, the subscript denotes the diffusion time step, where 0 corresponds to data and $1$ to pure noise. As discussed in Section 3.1, refining large homogeneous regions in an LR image is often unnecessary. To address this, we propose a mask-guided diffusion process, where the mask, often being sparse, is derived from the quadtree representation of the LR image. This approach reduces computational redundancy and reallocates model capacity toward refining detail-rich regions. In the following, we introduce the forward and reverse processes of QDM, along with the corresponding training objective.
Figure 2: Dual-Stream Architecture Overview. Our transformer-based model employs two complementary processing streams guided by a quadtree mask for input-adaptive computation. The upstream (top) extracts global context via Diffusion Transformer (DiT) blocks operating on patches with size $\times8$, while the downstream (bottom) refines patches with size $\times2$ exclusively in mask-selected regions. Mask $M$ dynamically controls computational overhead by routing computationally intensive processing only to detail-rich areas. Selected tokens are partitioned into parallelizable chunks, each processed independently through cross-attention blocks. Final predictions combine the coarse upstream outputs with the refined downstream results.
Forward Process. In the forward process, we gradually add noise to the HR image $Y_{0}$ conditioned on the LR image $X_{0}$. In contrast to vanilla diffusion models where the forward process adds noise everywhere, our approach only adds noise at active sites, i.e., detail-rich regions where the redundancy is low. In particular, we first obtain a mask $M$ of active sites via Algorithm 1, where entries of $M$ with value 1 indicate the active sites. Inspired by the work [yue2024resshift], we design the following transition distribution of the forward process that focuses solely on these active sites,
@@
Reverse Process. In the reverse process, we learn to perform denoising to recover the HR image. In particular, we have the following posterior distribution:
@@
where $p(X_{T}\mid Y_{0},M)={\mathcal{N}}\left(X_{T};Y_{0},\mathrm{\text{diag}}(\kappa^{2}M)\right)$ and $p_{\bm{\theta}}\left(X_{t-1}\mid X_{t},Y_{0},M\right)$ is the reverse transition probability. Following common practice in diffusion models [sohl2015deep, ho2020denoising, song2020score, yue2024resshift], we parameterize the reverse transition probability as:
@@
$\hat{\bm{\mu}}=\frac{\eta_{t-1}}{\eta_{t}}\odot X_{t}+\frac{\alpha_{t}}{\eta_{t}}\odot f_{\bm{\theta}}(X_{t},Y_{0},t),$ (4)
where $f_{\bm{\theta}}$ is a deep neural network with parameters $\bm{\theta}$, tailored to accurately estimate pixel values on active sites while making only minimal adjustments to the remaining regions. During the reverse process, the model iteratively denoises the active (detail-rich) regions, while applying minimal deterministic refinement to non-active (homogeneous) areas to preserve their structures. This procedure is illustrated in Figure 3. The training objective is to minimize the following mean squared error,
@@
We also adopted the non-uniform geometric schedule for the diffusion time $t$ as proposed in [yue2024resshift].
Extension to Latent Space. To reduce computational costs, recent studies [rombach2022high, peebles2023scalable, yue2024resshift] propose training diffusion models in the latent space of a pretrained variational autoencoder (VAE). In our framework, the encoder compresses images, reducing their spatial dimensions, which introduces a size mismatch between the compressed latent maps and the low-quality images. Fortunately, Alg. 1 can generate masks that align with the resolution of the latent maps. The spatial locations of active regions in the original image are preserved in the latent space due to the translation-equivariant properties of the convolutional encoder. Adapting our model to the latent space only requires replacing $X_{0}$ and $Y_{0}$ with their latent representations, along with the appropriately sized mask $M$.
@@ 3.3 Model Architecture
Fully transformer-based architectures [peebles2023scalable] have outperformed the conventional U-Net models [ronneberger2015u] in many tasks like image super-resolution. However, existing approaches are limited by their computationally expensive full-sequence attention mechanisms. This leads to substantial memory consumption and inefficient resource allocation, particularly in super-resolution tasks where only certain detail-rich regions necessitate complex processing, while others could be handled with simpler transformations.
@@
Upstream. The upstream branch utilizes a Diffusion Transformer (DiT) [peebles2023scalable] with a large patch size of 8 to extract global features from the input image. This design enables coarse refinements in homogeneous regions while reducing the token count by $64\times$ compared to pixel-level processing. By operating on coarse-grained patches, the model significantly lowers memory consumption and inference time while preserving global context. The large patch size expands the receptive field, improving efficiency in high-level semantic feature extraction. These global features are then propagated to downstream modules via cross-attention layers, enhancing spatial consistency in predictions.
Downstream. The downstream branch works in tandem with the upstream one, refining key regions identified by the quadtree mask $M$ (Alg. 1). Given an input of size $H\times W\times C$, the image is first divided into $2\times 2$ patches, forming a feature grid of dimensions $\frac{H}{2}\times\frac{W}{2}\times C$. Following the Swin Transformer [liu2021swin], this grid is further partitioned into non-overlapping $N\times N$ local windows, producing $\frac{HW}{4N^{2}}$ feature chunks of size $N^{2}\times C$. The quadtree mask $M$ (size $\frac{H}{2}\times\frac{W}{2}\times C$) selectively determines which chunks require computation: only windows overlapping with active sites (mask entries being $1$) are projected into a high-dimensional space. These selected features are then augmented with 2D positional embeddings to preserve spatial relationships.
Each feature chunk is processed through a series of attention-based operations. It first undergoes multi-head self-attention (MSA) to model local dependencies, followed by multi-head cross-attention (MCA) to incorporate global context from the upstream branch’s coarse features (dimension $C_{\text{up}}$, size $\frac{HW}{64}\times C_{\text{up}}$). The output is then passed through a two-layer feedforward network with GELU activation and time-conditioned via adaLN-Zero layers [peebles2023scalable]. Residual connections are applied after each MSA, MCA, and feedforward operation. To improve efficiency, the upstream’s global features are pre-projected into key-value pairs ($K,V$) and cached for reuse across all chunks during inference, minimizing redundant computations. The final projection restores features to their original dimensionality.
@@
where $\mathcal{L}_{\text{LPIPS}}$ [zhang2018unreasonable] measures perceptual similarity between prediction $f_{\bm{\theta}}(X_{t},Y_{0},t)$ and ground-truth $X_{0}$, with $\lambda_{3}$ balancing detail preservation against prediction accuracy.
Model Size. We apply a sequence of $N$ attention blocks for both upstream and downstream. Following DiT [peebles2023scalable], we use standard transformer configurations that jointly scale up the number of attention blocks, the hidden size, and the number of attention heads. Specifically, we use three configs, QDM-S, QDM-B and QDM-L, aligned in size with the DiT variants. Appendix A provides more details of these configurations.
Figure 5: Comparison of QDM, DiT [peebles2023scalable], U-Net [ronneberger2015u], and a Swin Attention-enhanced U-Net variant [liu2021swin] with high-resolution ($1024\times 1024$) inputs. The left panel shows peak memory usage, while the right panel depicts inference speed across evaluated models. All metrics were measured on an NVIDIA A100 80G GPU. For QDM, the evaluation was conducted under the most resource-intensive condition, using a full mask without adaptive computation and processing 64 chunks in parallel.